# Keeping in Place After the Storm—Emergency Assistance and Evictions

### v1

@@ 3.2.1 Hurricanes
We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance. We find that when FEMA assistance is made available, transactions volumes remain stable while default rates decrease comparatively, reflecting both a reduction in supply of credit when FEMA is not available but also an increase in precarity in making repayments. In particular, Figure 11 and Figure 12 shows that around weeks 6-10 post treatment we observe a spike in defaults and reduction in transaction volumes (possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 15-day loans). This also coincides around when evictions peak for non-FEMA receiving hurricanes in the previous section.
Specifically, in Table 13 we show that the volume of transactions reduces by 8.364 per week and defaults decrease overall by 1.458 per week on average per ZIP code. This reflects a 3.8% decrease in transactions and a 14% decrease in defaults per ZIP code per week relative to the baseline period. Comparatively, when FEMA rental assistance is not available available, transactions volumes decreased by 27.326 per week, a decline of 12.3%. While when FEMA assistance is available, defaults reduced by 1.941 per ZIP code per week relative to the baseline period, a decrease of 18.6% per week per ZIP code relative to the baseline period. And as Table 16 show, on the shorter interval of 7 weeks post hurricane event, this difference was even larger, an overall difference of 29.7% in weekly defaults between hurricane events in which FEMA rental assistance was and was not available.
@@
Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the transaction volume. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents estimates for ZIP codes that received FEMA aid, while panel Panel (b) shows estimates for ZIP codes that did not receive FEMA aid. The dependent variable is the default rates. Vertical lines represent 95% confidence intervals, constructed using standard errors clustered at the ZIP code level.
@@
Variable Control Group Treated Group
FEMA No FEMA FEMA No FEMA
Transaction Volume 209.930 283.549 226.306 202.092
Default 8.286 9.954 11.117 7.738
Table 15: Difference-in-Differences analysis of transaction volume and default rate outcomes at the ZIP code level, disaggregated by FEMA assistance. The analysis compares the impact of the hurricane event on ZIP codes with and without FEMA assistance over two weekly windows: from 5 weeks before to 7 weeks after the event ([-5 ; 7]) and from 5 weeks before to 12 weeks after the event ([-5 ; 12]). For the [-5 ; 7] window, ZIP codes with FEMA assistance experienced a small reduction in transaction volume (-2.998), while those without FEMA assistance saw a much larger decrease (-10.675∗∗∗). For default rates, FEMA-assisted ZIP codes experienced a slight reduction (-1.837∗∗∗), while those without FEMA saw an increase (1.253∗∗∗). For the longer [-5 ; 12] window, FEMA-assisted ZIP codes experienced a smaller reduction in transaction volume (-2.356), while non-FEMA-assisted ZIP codes saw a substantial decrease (-27.237∗∗∗). Default rates followed a similar pattern, with a decrease in FEMA-assisted ZIP codes (-1.941∗∗∗) and a small increase in non-FEMA ZIP codes (0.056).
Weekly Window [-5 ; 7]
@@
FEMA No FEMA FEMA No FEMA
Treated×$\times$×Post -2.998 -10.675∗∗∗ -1.837∗∗∗ 1.253∗∗∗
[2.750] [1.509] [0.548] [0.327]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 19517 6339 19517 6339
Weekly Window [-5 ; 12]
@@
FEMA No FEMA FEMA No FEMA
Treated×$\times$×Post -2.356 -27.237∗∗∗ -1.941∗∗∗ 0.056
[2.648] [2.878] [0.531] [0.827]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 27140 8772 27140 8772
Table 16: Triple Difference-in-Differences (DDD) analysis of transaction volume and default rate outcomes at the ZIP code level, examining the interaction between FEMA assistance and post-event outcomes. The table presents results for two weekly windows: [-5; 7] and [-5; 12]. For both windows, the treatment effect (Treated×$\times$×Post) is negative and statistically significant for transaction volume, indicating a post-event decline in economic activity. Conversely, the treatment effect is positive for default rates, suggesting an increase in loan defaults following the event. The interaction term (Treated×$\times$×Post×$\times$×FEMA) indicates that FEMA-assisted ZIP codes experienced a significant decrease in default rates in both windows, highlighting the mitigating effect of federal disaster assistance.
Weekly Window [-5 ; 7] [-5 ; 12]
Dependent Variable Transaction Default Transaction Default
Treated×$\times$×Post -10.675∗∗∗ 1.253∗∗∗ -27.237∗∗∗ 0.056
[1.509] [0.327] [2.877] [0.258]
Treated×$\times$×Post×$\times$×FEMA 7.677∗∗ -3.090∗∗∗ 24.881∗∗ -1.997∗∗∗
[3.088] [0.664] [4.142] [0.607]
Event-specific ZIP Code FE Y Y Y Y
Event-specific Time FE Y Y Y Y
Observations 25856 25856 35912 35912
@@ 4 Conclusion
Using granular data on evictions and loan-level data on payday loans, we provide evidence of disaster induced increase in evictions and reliance on high cost credit in the state of Florida following major storms and hurricanes. We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit. Overall, this suggests that major individual level household impacts such as access to housing are sensitive to the availability of emergency liquidity, supporting the need for such programs and the need to understand the availability of other forms of disaster insurance in a world where many locations will be subject to an increasing propensity of climate shock events.

### v2

@@ 3.2.1 Hurricanes
We further compare the usage of high cost credit during hurricane events with the availability of FEMA rental assistance. We find that both transaction volumes and default rates are lower post storm and that when FEMA assistance is made available, transactions volumes and default rates decrease by greater magnitude compared to when it is not made available, reflecting a reduction in demand of payday credit when FEMA is available but also a decrease in precarity in making repayments. In particular, Figure 10 we observe a decrease in defaults and reduction in transaction volumes (except a spike around 4,8 and 12 weeks, possibly reflecting a feature of payday loans that repeat customers often rollover their outstanding 30-day loans).
Specifically, in Table 12 we show that the volume of transactions reduces by 8.364 per week and a decrease in defaults by 1.458 per week on average per ZIP code. This reflects a 3.8% decrease in transactions per week relative to the baseline period and 14% decrease relative to the baseline period. Comparatively, when FEMA rental assistance is available, we find transactions volumes decrease by 12.597 per week relative to non-FEMA ZIP codes, a decline of 5.7% relative to the pre-event baseline. Furthermore, we find a decrease decrease in defaults by 3.318 per week, an approximately 31.8% relative to the pre-event baseline. While when ZIP codes receiving FEMA assistance are excluded, transaction volumes reduced by 8.538 per ZIP code per week relative to the baseline period, a decrease of 3.9% per week per ZIP code relative to the baseline period. Defaults, on the other hand, show a not statistically significant decrease of 0.433.
@@
Figure 11: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level for the treated only. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents transaction estimates for ZIP codes that received FEMA aid and were in the storm path, while panel Panel (b) shows default estimates for ZIP codes that received FEMA aid and were in the storm path. Standard errors are clustered at the ZIP code level.
@@
Period Weekly Weekly
Window [-5 ; 12] [-5 ; 12]
FEMA×$\times$×Post -12.597 -3.316∗∗∗
[7.130] [1.112]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 7366 7366
Table 13: Results for transaction volume and default rate outcomes at the ZIP code level for the payday loan data. The analysis evaluates the impact of hurricane event across treated groups and interacted with receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -12.597 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that received FEMA assistance, corresponding to approximately 5.7% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -3.316∗∗∗, indicating a significant reduction in the default rate, corresponding to approximately 31.8% lower default rates in treated ZIP codes that received FEMA assistance.
Figure 12: The figure provides the estimated coefficients from an event study examining the impact of hurricane exposure on weekly payday loan-related outcomes at the ZIP code level excluding ZIP codes that received FEMA assistance. The analysis spans a window from five weeks before to 12 weeks after the hurricane event, with the week immediately preceding the event (week -1) serving as the reference period. Panel (a) presents transaction volume estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance, while panel Panel (b) shows default estimates for ZIP codes relative to the control group excluding treated ZIP codes that received FEMA assistance. Standard errors are clustered at the ZIP code level.
@@
Period Weekly Weekly
Window [-5 ; 12] [-5 ; 12]
Treated×$\times$×Post -8.538∗∗∗ -0.433
[1.727] [0.348]
Event-specific ZIP Code FE Y Y
Event-specific Time FE Y Y
Observations 31686 31686
Table 14: Difference-in-differences esults for transaction volume and default rate outcomes at the ZIP code level for the payday loan data. The analysis evaluates the impact of hurricane event relative to the control group excluding treated ZIP codes receiving FEMA rental assistance over the weekly window: from 5 weeks before to 12 weeks after the event ([-5 ; 12]). The coefficient of -8.538 for transaction volume in the suggests a significant decrease in transaction volume in the treated ZIP codes that did not receive FEMA assistance relative to the control group, corresponding to approximately 3.9% lower transaction volume relative to pre-event levels. The coefficient for default rate in the same window is -0.433, indicating a non statistically significant reduction in the default rate when excluding treated ZIP codes that received FEMA assistance.
@@ 4 Conclusion
Using granular data on evictions and loan-level data on payday loans, we provide evidence of disaster induced increase in evictions and reliance on high cost credit in the state of Florida following major storms and hurricanes. We find that the existence of federal assistance helps mitigate the rate of evictions while also allowing households to avoid defaulting when relying on higher cost credit. Overall, this suggests that major household level impacts such as access to housing are sensitive to the availability of emergency liquidity, suggesting the utility for such programs and the need to better understand their availability in a world subject to an increasing propensity of climate shock events.